### v3

@@ 1.2 Technical overview
[…]
@@
To complement Theorem 1.3, we need to show how one can use the distribution (or equivalently, mixed-strategy for the $d$-player) provided by Theorem 1.3 to design a budget-feasible mechanism achieving a good approximation. Our mechanism (see Algorithm 1 in Section 4) builds upon the framework used by [BCGL17] (also later used by [NPS24]). We randomly partition the set of sellers into two groups $U_{1},U_{2}$, get an estimate of $\operatorname{OPT}$ from the sellers in $U_{1}$ who are then discarded, and work over the sellers in $U_{2}$. The set $S^{*}$ we use in Theorem 1.3 is obtained by computing a demand set over $U_{2}$ with appropriately chosen prices. This ensures that $v(S^{*})=\Omega(\operatorname{OPT})$. As discussed earlier, we sample a posted-payment vector $d$ from the distribution $\mathcal{D}$ in Theorem 1.3, which yields the set of elements $R=\{e\in S^{*}:c_{e}\leq d_{e}\}$. The value $d_{e}$ can be thought of as the threshold for seller $e$, in the sense that $e$ will not be picked into the solution if it declares a cost higher than $d_{e}$. So we often refer to the vector $d$ as a threshold vector for the elements in $S^{*}$. The resulting mechanism is truthful (due to Myerson’s characterization of truthfulness, Theorem 2.3) and has payments bounded by $d_{e}$, and since $d(U)\leq B$, we obtain budget-feasibility. If $c(S^{*})\leq\frac{B}{O(\log\log n)}$, the approximation factor follows from the guarantee in Theorem 1.3. But we also need to deal with the case where this does not hold. This presents some technical challenges, and here, we need to proceed significantly differently from the mechanisms of [NPS24, BCGL17]. [BCGL17] show in the analysis of their mechanism for subadditive valuations in the Bayesian setting, that when $c(S^{*})>\frac{B}{\alpha}$, the subadditive valuation $v$ is actually “$\alpha$-close to an XOS function”, and so one can use a mechanism for XOS valuations in this case. However, this requires one to have a suitable oracle for this XOS-approximation $v^{\mathsf{xos}}$ of $v$: either an XOS-oracle for $v^{\mathsf{xos}}$, or a demand oracle for $v^{\mathsf{xos}}$, as required by the mechanisms for XOS valuations in [BCGL17] and [NPS24] respectively. We only have a demand oracle for $v$, so this presents a significant challenge. In the mechanism of [NPS24] for subadditive valuations, $S^{*}$ is obtained using what they call a knapsack-cover oracle for $v$, which ensures that $c(S^{*})\leq B$, so they can always resort to Theorem 1.2. In our case, we do not have a knapsack-oracle for $v$, and we are using Theorem 1.3, so even if we had $c(S^{*})\leq B$, we still need to deal with the case where $c(S^{*})>B/O(\log\log n)$.
@@ 1.3 Related work
Budget-feasible mechanism design was introduced by Singer [Sin13], who also gave $O(1)$- randomized mechanisms for submodular valuations. This was followed by a sequence of work [BGG+22, BCGL17, ABM17, ABM16, AKS19, JT18, LMSZ21, GJLZ20, DPS11, CGL11] improving approximation ratios for budget feasible mechanisms under additive, submodular, XOS, or subadditive valuations. Variants of the problem were also considered, such as incorporating additional downward monotone set family constraints [ABM16, LMSZ21, NPS24, HHCT23], considering large market assumptions [AGN14], considering multi-unit generalizations [CC14, AKM+23, WZQ+19, KS22, QWC+20], multidimensional generalizations [NPS25], or considering more general budget constraints [NPS24]. Variants incorporating experimental design problems were also looked at [HIM14].
@@
The use of distributions with bounded marginals, as in $\bigl{(}\operatorname{LP}{(\kappa,S)}\bigr{)}$, has been considered before in the context of subadditive valuations [DKL20, BW23]. Importantly, [DKL20] use this LP to construct an $O(\log\log n)$-prophet inequality for online subadditive combinatorial auctions. The existence of a Nash equilibrium for our item-bidding zero-sum game that achieves payoff $\frac{1}{2}v(U)$ was implicit in the work of [BCGL17, NPS24]. A variant of this game is also considered in [FFGL15] with $m$ buyers who bid on $n$ items. Each buyer $i\in[m]$ has their own subadditive valuation over subsets of items. The utility of each buyer is the value of the set of items where she bids the highest. They showed that this game has price of anarchy at least $\frac{1}{2}$.
@@ 2 Preliminaries
For an element $e\in U$ and subset $S\subseteq U$, we use $v(e)$ as a shorthand for $v(\{e\})$, and $S+e$, $S-e$ as shorthands for $S\cup\{e\}$ and $S\setminus\{e\}$ respectively. For a set $S\subseteq U$ and a vector $p\in\mathbb{R}^{U}_{+}$, we use $p(S)$ as a shorthand to denote $\sum_{e\in S}p_{e}$. For $B\in\mathbb{R}_{+}$ and $S\subseteq U$, let $\Delta^{S}(B)=\{x\in\mathbb{R}^{S}_{+}:x(S)\leq B\}$. We use $\Delta(B)$ as a shorthand for $\Delta^{U}(B)$. For a cost vector $c\in\mathbb{R}^{U}$, we denote by $c_{-e}$ the vector in $\mathbb{R}^{U-e}$ obtained by dropping the coordinate corresponding to $e$. Moreover, for $c^{\prime}_{e}\in\mathbb{R}$, we denote by $(c^{\prime}_{e},c_{-e})$ the vector $x$ in $\mathbb{R}^{U}$ where $x_{e^{\prime}}=c_{e^{\prime}}$ for all $e^{\prime}\in U-e$, and $x_{e}=c^{\prime}_{e}$.
@@
A mechanism consists of an allocation rule $f$ along with a payment scheme $p$. In the setting of budget-feasible mechanism design, we are given as input the publicly-known information of the valuation function $v:2^{U}\to\mathbb{R}_{+}$, the budget $B\in\mathbb{R}_{+}$, and the seller’s reported costs $\{c_{e}\}_{e\in U}$. The allocation rule, or algorithm, $f$ outputs a subset $S\subseteq U$ of sellers whose items are purchased by the mechanism designer. We refer to the set $S$ returned by $f$ as the set of winners. The payment scheme $p$ then assigns to each seller a payment $p_{e}$. The utility of seller $e$, when her true private cost is $\overline{c}_{e}\geq 0$ is $u_{e}=p_{e}-\overline{c}_{e}$ if $e\in S$ and $u_{e}=p_{e}$ if $e\not\in S$. Each seller reports a cost that maximizes her own utility. Note that $f,\{p_{e}\}_{e\in U}$ are functions of $v,B$ and $c$. For ease of notation, we treat $v$ and $B$ as fixed and treat $f,\{p_{e}\}_{e\in U}$ as functions of $c$ only. Additionally, $u_{e}$ is a function of $v,B$, the reported costs $c$ and the true cost $\overline{c}_{e}$ of seller $e$. For ease of notation, we tread $v,B$ as fixed and refer to $u_{e}(\overline{c}_{e};c_{e},c_{-e})$ as the utility of seller $e$ when the reported costs are $c$ and the true cost of seller $e$ is $\overline{c}_{e}$. We seek a mechanism $\mathcal{M}=(f,p)$ satisfying the following properties.
@@ 3 Finding a good distribution over threshold vectors
We use the notation $\mathcal{D}[S^{*}]$ to emphasize that the distribution depends on the set $S^{*}\subseteq U$. We begin by proving the following lemma, which is a formal statement of Lemma 1.4. The lemma shows that the dual variables of $\bigl{(}\operatorname{LP}{(\kappa,S)}\bigr{)}$ can be used to find a pure strategy $d$ for the $d$-player that achieves a payoff of at least $v(S)-\operatorname{OPT_{LP}}(\kappa,S)$ against all vectors $c\in\mathbb{R}^{S}_{+}$ satisfying $c(S)\leq\kappa B$. Fix $S\subseteq U$ and $\kappa\in(0,1)$. Fix an optimal solution $(p,\mu)$ to $\bigl{(}\operatorname{DLP}({\kappa,S})\bigr{)}$. We define $d^{S,\kappa}$ to be the vector in +$\mathbb{R}^{U}_{+}$ given by $d^{S,\kappa}_{e}=p_{e}\cdot\frac{B}{p(S)}$ for $e\in S$ and $d^{S,\kappa}_{e}=0$ for $e\in U\setminus S$,
@@ 4 The budget feasible mechanism
We now use the result in Section 3 to design a randomized budget feasible mechanism for subadditive valuations that achieves an approximation factor of $O(\log\log n)$, thereby proving Theorem 1.1. As mentioned in Section 1.2, we first randomly partition the set of sellers into two groups $U_{1},U_{2}$, and get an estimate of $\operatorname{OPT}$ from the sellers in $U_{1}$ and discard them. We then work over the sellers in $U_{2}$. We compute a demand set over $U_{2}$, with appropriately chosen prices. This results in a set $S^{*}\subseteq U_{2}$ that has good value compared to $\operatorname{OPT}$. When $c(S^{*})\leq\frac{B}{O(\log\log n)}$, we can utilize Theorem 3.1 to prove a good approximation ratio by sampling a vector $d\sim\mathcal{D}[S^{*}]$ and return $R=\{e\in S^{*}:c_{e}\leq d_{e}\}$. However, we also need to deal with the case when $c(S^{*})>\frac{B}{O(\log\log n)}$. In this scenario, we will show that one can construct a single vector $d^{\prime}$ that provides suitable thresholds. We show that a suitable choice for $d^{\prime}$ is to set $d^{\prime}_{e}=q_{e}\cdot\frac{O(\log\log n)B}{\operatorname{OPT}}$, where $q\in\mathbb{R}^{S^{*}}_{+}$ is an optimal solution to the LP: $\max\ q(S^{*})$ subject to $q(S)\leq v(S^{*})-v(S^{*}\setminus S)$ for all $S\subseteq S^{*}$. We select the set of elements $A$ for which $c_{e}\leq d^{\prime}_{e}$, and prune this set to construct a subset $R^{\prime}\subseteq A$ satisfying $\Omega(\operatorname{OPT})-v(e^{*})\leq v(R^{\prime})\leq O(\operatorname{OPT})$. We show that constructing $R^{\prime}$ this way admits budget-feasible threshold payments. Returning either $R^{\prime}$ or $e^{*}$ then gives a good approximation, since if $v(e^{*})$ is small then $v(R^{\prime})\geq\Omega(\operatorname{OPT})-v(e^{*})$ would be large. One caveat here is that we cannot really branch on whether $c(S^{*})\leq\frac{B}{O(\log\log n)}$ as this causes problems with truthfulness. So instead, we return one of $R$, $R^{\prime}$, or $e^{*}$, each chosen with a suitable probability. The algorithm is described in detail below. We show that Algorithm 1 is a distribution over monotone algorithms, and so the corresponding payments can be obtained using Theorem 2.3.
@@
6Let $V_{1}$ be a $(2+\epsilon)$-approximation of $V_{1}^{*}$, the optimum value over $U_{2}$, obtained using Lemma 2.2. Let $\alpha=\frac{1}{4(2+\epsilon)+1}$ and $\beta=1-\alpha$.
@@
13Let $R^{\prime}$ be a maximal prefix of $A$ satisfying $q(R^{\prime})\leq\frac{1}{4}V_{1}$.
@@
The final condition above can be easily checked using a demand-oracle query with prices $q$ [^4] More precisely, we set the price of an element $e\in S^{*}$ to be $q_{e}$, and set the prices of elements outside of $S^{*}$ to be $\infty$, or some very large value. Then the demand-oracle query effectively amounts to a demand-oracle query over the ground set $S^{*}$. If the oracle returns some $T\subseteq S^{*}$ with $v(T)-q(T)>v(S^{*})-q(S^{*})$, then the LP constraint for $S^{*}\setminus T$ is violated, and otherwise $q$ is feasible. ∎
@@
[…]
@@
Note that the above discussion also shows that the threshold for a player $e$ can be computed in polytime. The threshold value $\tau^{\prime}_{e}$ for $e$ to belong to the set $S^{*}$ can be obtained using a demand oracle. If the set output is $R$, then the threshold for $e$ is $\min\{\tau^{\prime}_{e},d_{e}\}$. If the set output is $R^{\prime}$, then the threshold for $e$ to lie in $A$ is $\tau^{A}_{e}=\min\bigl{\{}\tau^{\prime}_{e},q_{e}\cdot\frac{4B}{V_{1}}\bigr{\}}$. Note that fixing $A$ also fixes the set $R^{\prime}$. Let $\overline{A}$ be the $A$-set obtained under input $(c^{\prime}_{e},c_{-e})$ when $e$ belongs to it, i.e., when $c^{\prime}_{e}<\tau^{A}_{e}$; let $\overline{R}$ be the corresponding $R^{\prime}$-set. Note $\overline{A}$ and $\overline{R}$ depend only on $c_{-e}$. One can infer that $\tau_{e}=\tau^{A}_{e}$, if $e\in\overline{R}$ and is 00 otherwise. ∎

### v5

@@ 1.2 Technical overview
Finally, to address (2), we show that the optimal solution to the dual of $\bigl(\operatorname{LP}{(\kappa,S)}\bigr)$ (see $\bigl(\operatorname{DLP}({\kappa,S})\bigr)$) can be used to design a pure strategy (that is, a vector of bids) for the $d$-player. This is captured by the following lemma.
@@
To complement Theorem 1.3, we need to show how one can use the distribution (or equivalently, mixed-strategy for the $d$-player) provided by Theorem 1.3 to design a budget-feasible mechanism achieving a good approximation. Our mechanism (see Algorithm 1 in Section 4) builds upon the framework used by [BCGL17] (also later used by [NPS24]). We randomly partition the set of sellers into two groups $U_{1},U_{2}$, get an estimate of $\operatorname{OPT}$ from the sellers in $U_{1}$ who are then discarded, and work over the sellers in $U_{2}$. The set $S^{*}$ we use in Theorem 1.3 is obtained by computing a demand set over $U_{2}$ with appropriately chosen prices. This ensures that $v(S^{*})=\Omega(\operatorname{OPT})$. As discussed earlier, we sample a posted-payment vector $d$ from the distribution $\mathcal{D}$ in Theorem 1.3, which yields the set of elements $R=\{e\in S^{*}:c_{e}\leq d_{e}\}$. The value $d_{e}$ can be thought of as the threshold for seller $e$, in the sense that $e$ will not be picked into the solution if it declares a cost higher than $d_{e}$. So we often refer to the vector $d$ as a threshold vector for the elements in $S^{*}$. The resulting mechanism is truthful (due to Myerson’s characterization of truthfulness, Theorem 2.3) and has payments bounded by $d_{e}$, and since $d(U)\leq B$, we obtain budget-feasibility. If $c(S^{*})\leq\frac{B}{O(\log\log n)}$, the approximation factor follows from the guarantee in Theorem 1.3. But we also need to deal with the case where this does not hold. This presents some technical challenges, and here, we need to proceed significantly differently from the mechanisms of [NPS24, BCGL17]. [BCGL17] show in the analysis of their mechanism for subadditive valuations in the Bayesian setting, that when $c(S^{*})>\frac{B}{\alpha}$, the subadditive valuation $v$ is actually $\alpha$-close to an XOS function”, and so one can use a mechanism for XOS valuations in this case. However, this requires one to have a suitable oracle for this XOS-approximation $v^{\mathsf{xos}}$ of $v$: either an XOS-oracle for $v^{\mathsf{xos}}$, or a demand oracle for $v^{\mathsf{xos}}$, as required by the mechanisms for XOS valuations in [BCGL17] and [NPS24] respectively. We only have a demand oracle for $v$, so this presents a significant challenge. In the mechanism of [NPS24] for subadditive valuations, $S^{*}$ is obtained using what they call a knapsack-cover oracle for $v$, which ensures that $c(S^{*})\leq B$, so they can always resort to Theorem 1.2. In our case, we do not have a knapsack-oracle for $v$, and we are using Theorem 1.3, so even if we had $c(S^{*})\leq B$, we still need to deal with the case where $c(S^{*})>B/O(\log\log n)$.
@@ 1.3 Related work
Budget-feasible mechanism design was introduced by Singer [Sin13], who also gave $O(1)$-factor randomized mechanisms for submodular valuations. This was followed by a sequence of work [BGG+22, BCGL17, ABM17, ABM16, AKS19, JT18, LMSZ21, GJLZ20, DPS11, CGL11] improving approximation ratios for budget feasible mechanisms under additive, submodular, XOS, or subadditive valuations. Variants of the problem were also considered, such as incorporating additional downward monotone set family constraints [ABM16, LMSZ21, NPS24, HHCT23], considering large market assumptions [AGN14], considering multi-unit generalizations [CC14, AKM+23, WZQ+19, KS22, QWC+20], multidimensional generalizations [NPS25], or considering more general budget constraints [NPS24]. Variants incorporating experimental design problems were also looked at [HIM14].
@@
The use of distributions with bounded marginals, as in $\bigl(\operatorname{LP}{(\kappa,S)}\bigr)$, has been considered before in the context of subadditive valuations [DKL20, BW23]. Importantly, [DKL20] use this LP to construct an $O(\log\log n)$-prophet inequality for online subadditive combinatorial auctions. The existence of a Nash equilibrium for our item-bidding zero-sum game that achieves payoff $\frac{1}{2}v(U)$ was implicit in the work of [BCGL17, NPS24]. Variants of our item-bidding game, in the context of first-price and second-price auctions over subadditive valuations, have been analyzed in [BR11, HKMN11, FFGL15]. Our item-bidding game can also be seen as a generalization of the Colonel Blotto game to subadditive valuations [Bor53]. The Colonel Blotto game has attracted recent attention in theoretical computer science [Rob06, VLS18, BAEJ20, BDD+22].
@@ 2 Preliminaries
For an element $e\in U$ and subset $S\subseteq U$, we use $v(e)$ as a shorthand for $v(\{e\})$, and $S+e$, SeS-e as shorthands for $S\cup\{e\}$ and $S\setminus\{e\}$ respectively. For a set $S\subseteq U$ and a vector $p\in\mathbb{R}^{U}_{+}$, we use $p(S)$ as a shorthand to denote $\sum_{e\in S}p_{e}$. For $B\in\mathbb{R}_{+}$ and $S\subseteq U$, let $\Delta^{S}(B)=\{x\in\mathbb{R}^{S}_{+}:x(S)\leq B\}$. We use $\Delta(B)$ as a shorthand for $\Delta^{U}(B)$. For a cost vector $c\in\mathbb{R}^{U}$, we denote by $c_{-e}$ the vector in $\mathbb{R}^{U-e}$ obtained by dropping the coordinate corresponding to $e$. Moreover, for $c^{\prime}_{e}\in\mathbb{R}$, we denote by $(c^{\prime}_{e},c_{-e})$ the vector $x$ in $\mathbb{R}^{U}$ where $x_{e^{\prime}}=c_{e^{\prime}}$ for all $e^{\prime}\in U-e$, and $x_{e}=c^{\prime}_{e}$.
@@
A mechanism consists of an allocation rule $f$ along with a payment scheme $p$. In the setting of budget-feasible mechanism design, we are given as input the publicly-known information of the valuation function $v:2^{U}\to\mathbb{R}_{+}$, the budget $B\in\mathbb{R}_{+}$, and the seller’s reported costs $\{c_{e}\}_{e\in U}$. The allocation rule, or algorithm, $f$ outputs a subset $S\subseteq U$ of sellers whose items are purchased by the mechanism designer. We refer to the set $S$ returned by $f$ as the set of winners. The payment scheme $p$ then assigns to each seller a payment $p_{e}$. The utility of seller $e$, when her true private cost is $\overline{c}_{e}\geq 0$ is $u_{e}=p_{e}-\overline{c}_{e}$ if $e\in S$ and $u_{e}=p_{e}$ if $e\not\in S$. Each seller reports a cost that maximizes her own utility. Note that $f,\{p_{e}\}_{e\in U}$ are functions of $v,B$ and $c$. For ease of notation, we treat $v$ and $B$ as fixed and treat $f,\{p_{e}\}_{e\in U}$ as functions of $c$ only. Additionally, $u_{e}$ is a function of $v,B$, the reported costs $c$ and the true cost $\overline{c}_{e}$ of seller $e$. For ease of notation, we treat $v,B$ as fixed and refer to $u_{e}(\overline{c}_{e};c_{e},c_{-e})$ as the utility of seller $e$ when the reported costs are $c$ and the true cost of seller $e$ is $\overline{c}_{e}$. We seek a mechanism $\mathcal{M}=(f,p)$ satisfying the following properties.
@@ 3 Finding a good distribution over threshold vectors
We use the notation $\mathcal{D}[S^{*}]$ to emphasize that the distribution depends on the set $S^{*}\subseteq U$. We begin by proving the following lemma, which is a formal statement of Lemma 1.4. The lemma shows that the dual variables of $\bigl(\operatorname{LP}{(\kappa,S)}\bigr)$ can be used to find a pure strategy $d$ for the $d$-player that achieves a payoff of at least $v(S)-\operatorname{OPT_{LP}}(\kappa,S)$ against all vectors $c\in\mathbb{R}^{S}_{+}$ satisfying $c(S)\leq\kappa B$. Fix $S\subseteq U$ and $\kappa\in(0,1)$. Fix an optimal solution $(p,\mu)$ to $\bigl(\operatorname{DLP}({\kappa,S})\bigr)$. We define $d^{S,\kappa}$ to be the vector in $\mathbb{R}^{U}_{+}$ given by $d^{S,\kappa}_{e}=p_{e}\cdot\frac{B}{p(S)}$ for $e\in S$ and $d^{S,\kappa}_{e}=0$ for $e\in U\setminus S$,
@@ 4 The budget feasible mechanism
We now use the result in Section 3 to design a randomized budget feasible mechanism for subadditive valuations that achieves an approximation factor of $O(\log\log n)$, thereby proving Theorem 1.1. As mentioned in Section 1.2, we first randomly partition the set of sellers into two groups $U_{1},U_{2}$, and get an estimate of $\operatorname{OPT}$ from the sellers in $U_{1}$ and discard them. We then work over the sellers in $U_{2}$. We compute a demand set over $U_{2}$, with appropriately chosen prices. This results in a set $S^{*}\subseteq U_{2}$ that has good value compared to $\operatorname{OPT}$. When $c(S^{*})\leq\frac{B}{O(\log\log n)}$, we can utilize Theorem 3.1 to prove a good approximation ratio by sampling a vector $d\sim\mathcal{D}[S^{*}]$ and return $R=\{e\in S^{*}:c_{e}\leq d_{e}\}$. However, we also need to deal with the case when $c(S^{*})>\frac{B}{O(\log\log n)}$. In this scenario, we will show that one can construct a single vector $d^{\prime}$ that provides suitable thresholds. We show that a suitable choice for $d^{\prime}$ is to set $d^{\prime}_{e}$ to be a suitable scaling of $q_{e}$, where $q\in\mathbb{R}^{S^{*}}_{+}$ is an optimal solution to the LP: $\max\ q(S^{*})$ subject to $q(S)\leq v(S^{*})-v(S^{*}\setminus S)$ for all $S\subseteq S^{*}$. We select the set of elements $A$ for which $c_{e}\leq d^{\prime}_{e}$, and prune this set to construct a subset $R^{\prime}\subseteq A$ satisfying $\Omega(\operatorname{OPT})-v(e^{*})\leq v(R^{\prime})\leq O(\operatorname{OPT})$. We show that constructing $R^{\prime}$ this way admits budget-feasible threshold payments. Returning either $R^{\prime}$ or $e^{*}$ then gives a good approximation, since if $v(e^{*})$ is small then $v(R^{\prime})\geq\Omega(\operatorname{OPT})-v(e^{*})$ would be large. One caveat here is that we cannot really branch on whether $c(S^{*})\leq\frac{B}{O(\log\log n)}$ as this causes problems with truthfulness. So instead, we return one of $R$, $R^{\prime}$, or $e^{*}$, each chosen with a suitable probability. The algorithm is described in detail below. We show that Algorithm 1 is a distribution over monotone algorithms, and so the corresponding payments can be obtained using Theorem 2.3.
@@
6Let $V_{1}$ be a $(2+\epsilon)$-approximation of $V_{1}^{*}$, the optimum value over $U_{2}$, obtained using Lemma $2.$ Let $\alpha=\frac{1}{4(2+\epsilon)+1}$ and $\beta=1-\alpha$.
@@
13Let $R^{\prime}$ be a maximal prefix of AA satisfying $q(R^{\prime})\leq\frac{1}{4}V_{1}$.
@@
The final condition above can be easily checked using a demand-oracle query with prices qq. [^4] More precisely, we set the price of an element $e\in S^{*}$ to be $q_{e}$, and set the prices of elements outside of $S^{*}$ to be $\infty$, or some very large value. Then the demand-oracle query effectively amounts to a demand-oracle query over the ground set $S^{*}$. If the oracle returns some $T\subseteq S^{*}$ with $v(T)-q(T)>v(S^{*})-q(S^{*})$, then the LP constraint for $S^{*}\setminus T$ is violated, and otherwise $q$ is feasible. ∎
@@
$\begin{split}0.2\cdot v(e^{*})+\frac{1}{4}\cdot 0.8\beta\cdot\tfrac{\alpha}{\beta}\cdot\min\Bigl\{\operatorname{OPT}-&\tfrac{v(e^{*})}{\alpha},\,\tfrac{\operatorname{OPT}-v(e^{*})}{64\log\log n}\Bigr\}\geq\min\Bigl\{0.2\alpha\cdot\operatorname{OPT},\,0.2\alpha\cdot\tfrac{\operatorname{OPT}}{64\log\log n}\Bigr\}\\ &=0.2\alpha\cdot\frac{\operatorname{OPT}}{64\log\log n}=\frac{\operatorname{OPT}}{(2880+1280\epsilon)\log\log n}.\end{split}$
@@
Note that the above discussion also shows that the threshold for a player $e$ can be computed in polytime. The threshold value $\tau^{\prime}_{e}$ for $e$ to belong to the set $S^{*}$ can be obtained using a demand oracle. If the set output is $R$, then the threshold for $e$ is $\min\{\tau^{\prime}_{e},d_{e}\}$. If the set output is $R^{\prime}$, then the threshold for $e$ to lie in $A$ is $\tau^{A}_{e}=\min\bigl\{\tau^{\prime}_{e},q_{e}\cdot\frac{4B}{V_{1}}\bigr\}$. Note that fixing $A$ also fixes the set $R^{\prime}$. Let $\overline{A}$ be the $A$-set obtained under input $(c^{\prime}_{e},c_{-e})$ when $e$ belongs to it, i.e., when $c^{\prime}_{e}<\tau^{A}_{e}$; let $\overline{R}$ be the corresponding $R^{\prime}$-set. Note $\overline{A}$ and $\overline{R}$ depend only on $c_{-e}$. One can infer that $\tau_{e}=\tau^{A}_{e}$, if $e\in\overline{R}$ and is 0 otherwise. ∎